OpenStudy (anonymous):

Translate to an inequality. Use the variable x. The number of people in the chess club is less than or equal to 23

OpenStudy (anonymous):

First of all, do you know what an inequality is?

OpenStudy (anonymous):

nequality is a relation that holds between two values

OpenStudy (anonymous):

Great, so when we are given the phrase 'less than or equal to' what two symbols can we use?

OpenStudy (anonymous):

@jlongSwag27 You there?

OpenStudy (anonymous):

>-

OpenStudy (anonymous):

@BlossomCake

OpenStudy (anonymous):

\[You~mean~\le~or~\ge~?\]

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

Okay, so do you know what to do now?

OpenStudy (anonymous):

no.....

OpenStudy (anonymous):

The sentence reads: The number of people in the chess club is less than or equal to 23. So, let's use the variable 'x' to stand for 'the number of people in the chess club' okay?

OpenStudy (anonymous):

x>23

OpenStudy (anonymous):

@BlossomCake

OpenStudy (anonymous):

with the live under >

OpenStudy (anonymous):

Actually, I think it might be: \[x~\le~23\]

OpenStudy (anonymous):

line*

OpenStudy (anonymous):

no thats greater than.. it says less than

OpenStudy (anonymous):

youre right my bad lol

OpenStudy (anonymous):

\[x~\le~23\] Reads: x is less than or equal to 23, right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

So, there's your inequality!

OpenStudy (anonymous):

\[\huge \dddot \smile\]